The formula s = 16t^2 is used to approximate the distance s, in feet, that an object falls freely (from rest) in t seconds. Use this formula to solve the problem. (Round answer to the nearest tenth.)
How long would it take an object to fall freely from a bridge 725 feet above the water

Answer by stanbon(75887)

The formula s = 16t^2 is used to approximate the distance s, in feet, that an object falls freely (from rest) in t seconds. Use this formula to solve the problem.
------
How long would it take an object to fall freely from a bridge 725 feet above the water
725 =16t^2
t^2 =725/16
t = sqrt(725)/4 seconds
=============================
Cheers,
Stan H.